# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ Abstract
Adequately generating and evaluating prediction models based on supervised machine learning (ML) is often challenging, especially for less experienced users in applied research areas. Special attention is required in settings where the model generation process involves hyperparameter tuning, i.e. data-driven optimization of different types of hyperparameters to improve the predictive performance of the resulting model. Discussions about tuning typically focus on the hyperparameters of the ML algorithm (e.g., the minimum number of observations in each terminal node for a tree-based algorithm). In this context, it is often neglected that hyperparameters also exist for the preprocessing steps that are applied to the data before it is provided to the algorithm (e.g., how to handle missing feature values in the data). As a consequence, users experimenting with different preprocessing options to improve model performance may be unaware that this constitutes a form of hyperparameter tuning - albeit informal and unsystematic - and thus may fail to report or account for this optimization. To illuminate this issue, this paper reviews and empirically illustrates different procedures for generating and evaluating prediction models, explicitly addressing the different ways algorithm and preprocessing hyperparameters are typically handled by applied ML users. By highlighting potential pitfalls, especially those that may lead to exaggerated performance claims, this review aims to further improve the quality of predictive modeling in ML applications.
@@ 1 Introduction
Many applied research areas have recently seen an increase in the development of prediction models based on supervised machine learning (ML) algorithms. However, after initially generating widespread enthusiasm – partly due to the availability of user-friendly software that enables model development without requiring extensive expertise – ML-based prediction models are now undergoing critical reexamination (Ball,, 2023; Kapoor and Narayanan,, 2023; Pfob et al.,, 2022). Among other concerns, such as insufficient reporting of relevant aspects of the model development process, it has been found that the claimed predictive performance of many models is considerably exaggerated (Kapoor and Narayanan,, 2023; Dhiman et al., 2022a, ; Dhiman et al., 2022b, ; Andaur Navarro et al.,, 2021). While some of the pitfalls leading to such optimistically biased performance claims (e.g., using the exact same observations for model generation and evaluation) typically occur only among very inexperienced applied ML users and are well known within the ML research community, others arise more subtly (Domingos,, 2012; Kapoor and Narayanan,, 2023; Poldrack et al.,, 2020; Hofman et al.,, 2023).
@@
This is particularly true when the model generation process involves data-driven hyperparameter optimization, which is also referred to as hyperparameter tuning and is commonly employed in ML applications. The most prominent type of hyperparameters (HPs) are those associated with the learning algorithm, which specify its configuration (e.g., the minimum number of observations in each terminal node for tree-based algorithms). If selected by an adequate (and ideally automated) tuning procedure, HPs can substantially enhance the performance of the resulting prediction model. However, HP tuning also complicates model evaluation, as common procedures such as simple $k$-fold cross-validation no longer guarantee an unbiased assessment (Bischl et al.,, 2023; Hosseini et al.,, 2020).
@@
An additional challenge comes from the fact that, beyond algorithm HPs, there are also preprocessing HPs, which specify the steps applied to the data before it is fed into the learning algorithm (e.g., selecting the set of features for prediction or determining how missing feature values are handled; Binder and Pfisterer,, 2024; Bischl et al.,, 2023). While the tuning of algorithm HPs is rightfully considered important for model performance, the relevance of tuning preprocessing HPs should not be overlooked. Preprocessing steps can make or break a model’s predictive performance, and solely relying on user expertise to specify these steps (which is the alternative to tuning) is often impractical and may result in arbitrary decisions (Kuhn and Johnson,, 2013). Despite this, reports of tuning preprocessing HPs aside from feature selection are relatively rare. This could be because integrating preprocessing HPs into automated tuning workflows typically requires advanced programming expertise, which not all applied ML users have, or because this possibility is not widely recognized. Importantly, the limited use of automated tuning procedures for preprocessing HPs does not mean that these HPs are not being tuned at all. In fact, it appears fairly common for applied ML users to experiment informally with different preprocessing options (Hosseini et al.,, 2020; Hofman et al.,, 2023; Lones,, 2024), often without realizing that this constitutes a form of (manual) HP tuning. If this type of tuning is indeed conducted subconsciously, it will also remain unaccounted for during model evaluation, thereby increasing the risk of drawing overly optimistic conclusions about the model’s performance.
@@
To avoid such issues, it is essential to educate users in applied settings about the different types of HPs, the different forms of HP tuning, and how tuning can impact both the true and estimated performance of prediction models. Although valuable literature already exists describing the concept of HP tuning and various automated procedures (e.g., Bischl et al.,, 2023; Bartz et al.,, 2023; Feurer and Hutter,, 2019), this research primarily adopts the perspective of ML methods researchers who are concerned with evaluating the overall performance of ML algorithms used to generate prediction models. This focus does not align with the perspective of applied ML users, who are more interested in the performance of a specific prediction model. Although this literature is still useful for them – since the general principles described there essentially hold for all types of audiences – applied ML users additionally need specific guidance for developing their “final model” (a notion that does not exist in the methodological context). Moreover, they may find it challenging to extract the relevant insights from literature aimed at a different audience with partly different needs. In contrast, literature explicitly directed toward applied ML users tends to either focus on general guidelines for ML-based predictive modeling, lacking detailed coverage of HP tuning (e.g., Kuhn and Johnson,, 2013; Pfob et al.,, 2022; Lones,, 2024; Kapoor et al.,, 2024; Poldrack et al.,, 2020; Collins et al., 2024a, ; Van Royen et al.,, 2023), or addresses HP tuning only within specific research areas (e.g., Hosseini et al.,, 2020; Dunias et al.,, 2024). Additionally, much of the existing HP tuning literature does not consider preprocessing HPs. Exceptions include the review by Bischl et al., (2023), which, however, touches on this topic only briefly. This lack of detail is reasonable, given that preprocessing HPs can, in principle, be tuned using the same automated procedures as algorithm HPs. However, this perspective overlooks that preprocessing HPs are often tuned manually in applied settings, which carries implications different from those associated with automated tuning.
@@ 2.1 Terminology and notation
The following terminology and notation is adapted from Bischl et al., (2023). Let $\mathcal{D}_{\text{train}}$ be a labeled data set with $n_{\text{train}}$ observations. Accordingly, each observation $i$ ($i=1,\dots,n_{\text{train}}$) consists of an outcome $y^{(i)}$ (i.e. the variable to be predicted, also referred to as label or target) and a $p$-dimensional feature vector $\boldsymbol{x}^{(i)}$ (i.e. the $p$ variables used to predict $y^{(i)}$, also referred to as predictors), where $y^{(i)}$ and $\boldsymbol{x}^{(i)}$ can take any value from the outcome space $\mathcal{Y}$ and feature space $\mathcal{X}$, respectively. Two common types of prediction problems are regression, for which $y^{(i)}$ can be any real number (i.e. $\mathcal{Y}=\mathds{R}$), and classification, for which $y^{(i)}$ can be one of $g$ classes (i.e. $\mathcal{Y}$ is finite and categorical with $|\mathcal{Y}|=g$). We assume that the observations in $\mathcal{D}_{\text{train}}$ are independent and have been sampled from the same (unknown) probability distribution $\mathds{P}_{xy}$.
@@
The general aim of supervised ML is to “learn” a model from the data set $\mathcal{D}_{\text{train}}$ that is able to predict the outcome values of new observations. Essentially, a prediction model is a function $\hat{f}:\mathcal{X}\rightarrow\mathds{R}^{g}$ that maps any observed feature vector $\boldsymbol{x}$ to a prediction vector $\hat{f}(\boldsymbol{x})$ in $\mathds{R}^{g}$. The prediction vector $\hat{f}(\boldsymbol{x})$ either directly corresponds to the predicted outcome value (e.g., for regression, where $g=1$), or can be transformed accordingly (e.g., for classification, where $\hat{f}(\boldsymbol{x})$ corresponds to predicted probabilities for each class and the predicted class could be the class with the highest probability). The prediction model results from a learning pipeline $\mathcal{I}$, which uses the data set $\mathcal{D}_{\text{train}}$ to find the function $\hat{f}$ that yields the best predictions for the true outcome values in $\mathcal{D}_{\text{train}}$. To stress that a prediction model $\hat{f}$ is based on learning pipeline $\mathcal{I}$ and data set $\mathcal{D}_{\text{train}}$, we write $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$. The prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can usually be parameterized, meaning that it is defined by a set of parameters $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ (simply denoted as $\boldsymbol{\hat{\theta}}$ when data set and learning pipeline are clear from context and $\boldsymbol{\theta}$ when referring to the parameters prior to estimation).
@@
There are two key processes associated with $\mathcal{I}$ and $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$, which we will explore in more detail throughout the paper: (i) the training process, in which the learning pipeline $\mathcal{I}$ is applied to $\mathcal{D}_{\text{train}}$ and estimates the parameters $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ and thus the prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$, and (ii) the prediction process, in which $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ is used to make predictions for an observation (whether from $\mathcal{D}_{\text{train}}$ or from a new data set) with feature vector $\boldsymbol{x}$, resulting in $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}(\boldsymbol{x})$. Note that to make predictions on a new data set, the outcome does not need to be observed (it would only be necessary for evaluating those predictions). The training and prediction process serve as the foundation for more complex processes related to the development of prediction models, which we will address in Section 2.4.
@@ 2.2 Learning pipeline
Each learning pipeline $\mathcal{I}$ contains a learning algorithm as a central component but can also include several preprocessing steps that are performed before the algorithm is applied to the data. Since preprocessing steps are a particular focus of this paper, we use the term “learning pipeline” instead of the more common term “learner” to emphasize that $\mathcal{I}$ can consist of several components. Note that for now, we consider all components of $\mathcal{I}$ as fixed, but will discuss the case in which they can be modified in Section 2.3.
@@ 2.2.1 Learning algorithm
The choice of learning algorithm usually depends on the specific prediction problem. For example, if the desired prediction model is a decision tree (which is the case for the real-world prediction problem considered in Section 5), a possible algorithm choice is the well-known Classification and Regression Tree algorithm (CART), which partitions the feature space $\mathcal{X}$ by a sequence of binary splits into terminal nodes and assigns a prediction value to each terminal node (Breiman et al.,, 1984). In this case, the parameters of the learning algorithm contained in $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ are the splitting rules that generate the tree structure (i.e. which features are used with which threshold value) and the prediction values at each terminal node. The learning algorithm can also consist of multiple individual algorithms that are combined into one overall algorithm (e.g., random forests). These types of algorithms are referred to as ensemble methods, but will not be discussed further in this paper. In general, the choice of algorithm has a large impact on the hypothesis space of the learning pipeline, i.e. the set of prediction models the learning pipeline can generate. For example, selecting a standard linear regression as algorithm (with $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ containing the regression coefficients) would imply that the corresponding learning pipeline would not be able to learn prediction models that do not correspond to linear combinations of the features (e.g., polynomials).
@@ 2.2.2 Preprocessing
While a data set can, in theory, be fed directly into the algorithm (i.e. the algorithm is the only component of the learning pipeline), it typically undergoes some modification first. This process can be referred to as data preprocessing and encompasses all the steps taken to transform the data set from its rawest available form into the final form provided as input to the learning algorithm (Kapoor et al.,, 2024). Data preprocessing steps are usually performed to improve the performance of the resulting prediction model, to enable the data to be (better) handled by the learning algorithm (Thomas,, 2024), or to improve the interpretability of the resulting prediction model. To better illustrate different the characteristics of preprocessing steps and their implications on the training and prediction process, we consider a simple learning pipeline as an example, which is also depicted in Figure 1 (middle panel).
@@
Based on this example learning pipeline, we can make a first distinction between preprocessing steps. This distinction concerns whether the steps have parameters estimated from $\mathcal{D}_{\text{train}}$ (with these parameters included in $\boldsymbol{\theta}$) or whether they are parameterless and are carried out independently for each observation (Kapoor et al.,, 2024; Binder and Pfisterer,, 2024). In the example, the replacement of missing feature values is a parameterized preprocessing step, as it involves the parameter $\theta_{impute}$, representing the mean of all non-missing values estimated from $\mathcal{D}_{\text{train}}$. In contrast, the log-transformation of features does not involve any parameters. Other examples of preprocessing steps with parameters include centering or scaling of features, where parameters such as the mean or standard deviation are estimated from $\mathcal{D}_{\text{train}}$. On the other hand, creating a new feature by summing multiple features serves as another example of a parameterless preprocessing step.
@@
We first consider preprocessing steps that affect only the feature distribution. These comprise all preprocessing steps mentioned above, including those in the example learning pipeline. Additional examples are dimensionality reduction techniques (e.g., principal component analysis), feature selection, or data cleaning steps that do not alter the outcome distribution (e.g., correction of errors in features) (Kuhn and Johnson,, 2013; Thomas,, 2024). Preprocessing steps of this type must be applied not only during training but also during prediction, in the same sequence as in the learning pipeline. This ensures that the model produced by the learning algorithm receives the data in the same format during prediction as it did during training, preserving the validity of the model (Binder and Pfisterer,, 2024). This requirement implies that these steps are not only components of the learning pipeline $\mathcal{I}$, but also part of the resulting prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$. Consequently, if a learning pipeline $\mathcal{I}$ includes $h$ preprocessing steps that only affect the feature distribution, the prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ is not a single function but a function composition of $h+1$ functions (omitting $\mathcal{D}_{\text{train}}$ and $\mathcal{I}$ for simplicity of notation):
@@
where $\hat{f}_{h+1}$ corresponds to the model resulting from the learning algorithm, and $\hat{f}_{h},\dots,\hat{f}_{1}$ reflect the $h$ preprocessing steps. Accordingly, a more accurate name for a prediction model would be prediction model pipeline, but for brevity, we will continue to use the former. Returning to the example learning pipeline, the resulting prediction model is a composition of three functions, $\hat{f}_{3}(\hat{f}_{2}(\hat{f}_{1}(\boldsymbol{x})))$, where $\hat{f}_{1}$, $\hat{f}_{2}$, and $\hat{f}_{3}$ correspond to the imputation step, the log-transformation step, and the decision tree model, respectively. When making a prediction for one or more observations, all three functions must be applied (see Figure 1, right panel). Importantly, if any functions constituting the prediction model are omitted during the prediction process, or if any preprocessing or algorithm parameters are re-estimated on a new data set for which predictions are to be made, the validity of the prediction model may be compromised. However, in practice, this pitfall is often unavoidable for users who wish to apply a model but were not involved in its development, as studies introducing new prediction models frequently fail to report the preprocessing steps performed prior to applying the learning algorithm (Kapoor et al.,, 2024).
@@
In contrast to preprocessing steps that only affect the feature distribution, preprocessing steps that modify the outcome distribution are not necessarily applied during prediction. Here, we must distinguish between steps aimed at improving compatibility with the learning algorithm and those intended to alter the scope or interpretation of the prediction model. An example of the first type is (invertible) transformations applied to the outcome during training, such as a log-transformation to reduce skewness. To ensure predictions are returned on the correct scale, these transformations must be reversed during prediction (Thomas,, 2024). For instance, if the outcome was log-transformed during training, the model will output $\mathrm{log}(\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}(\boldsymbol{x}))$, which must then be exponentiated to restore the prediction to its original scale. Note that some other compatibility-focused steps are not applied at all during prediction. In the context of classification problems, this includes class-balancing steps such as oversampling, where observations from the least prevalent class are randomly resampled to overcome class imbalance effects during the training process (see, e.g., Kuhn and Johnson,, 2013 for more details). In the notation of the prediction model as a function composition introduced above, preprocessing steps that are applied only in their inverted form or not at all during prediction are represented as inversion function or identity function, respectively.
@@
In contrast, preprocessing steps that modify the outcome to alter the scope or interpretation of the prediction model should be consistently applied during prediction. For example, if a continuous outcome is discretized to convert a regression problem into a classification problem (Hofman et al.,, 2023), this (irreversible) transformation must also be applied to the true outcome during prediction in order to enable a meaningful comparison between the predictions and the actual outcome values. Such transformations of the outcome are not part of the prediction model itself (which maps $\boldsymbol{x}$ to predictions, not $y$), but must be performed alongside the prediction process. Moreover, since the outcome values are generally unknown when making predictions for observations from a new data set that does not correspond to $\mathcal{D}_{\text{train}}$, these transformations are typically not actual steps executed when making predictions but instead determine how the predictions are interpreted.

### v2

@@ Abstract
Adequately generating and evaluating prediction models based on supervised machine learning (ML) is often challenging, especially for less experienced users in applied research areas. Special attention is required in settings where the model generation process involves hyperparameter tuning, i.e. data-driven optimization of different types of hyperparameters to improve the predictive performance of the resulting model. Discussions about tuning typically focus on the hyperparameters of the ML algorithm (e.g., the minimum number of observations in each terminal node for a tree-based algorithm). In this context, it is often neglected that hyperparameters also exist for the preprocessing steps that are applied to the data before it is provided to the algorithm (e.g., how to handle missing feature values in the data). As a consequence, users experimenting with different preprocessing options to improve model performance may be unaware that this constitutes a form of hyperparameter tuning, albeit informal and unsystematic, and thus may fail to report or account for this optimization. To illuminate this issue, this paper reviews and empirically illustrates different procedures for generating and evaluating prediction models, explicitly addressing the different ways algorithm and preprocessing hyperparameters are typically handled by applied ML users. By highlighting potential pitfalls, especially those that may lead to exaggerated performance claims, this review aims to further improve the quality of predictive modeling in ML applications.
Keywords: predictive modeling, machine learning, preprocessing, hyperparameter optimization, tuning
@@ 1 Introduction
Many applied research areas have recently seen an increase in the development of prediction models based on supervised machine learning (ML) algorithms. However, after initially generating widespread enthusiasm—partly due to the availability of user-friendly software that enables model development without requiring extensive expertise—ML-based prediction models are now undergoing critical reexamination [75, 108, 121]. Among other concerns, such as insufficient reporting of relevant aspects of the model development process, it has been found that the claimed predictive performance of many models is considerably exaggerated [108, 88, 89, 73]. While some of the pitfalls leading to such optimistically biased performance claims (e.g., using the exact same observations for model generation and evaluation) typically occur only among very inexperienced applied ML users and are well known within the ML research community, others arise more subtly [90, 108, 122, 100].
@@
This is particularly true when the model generation process involves data-driven hyperparameter optimization, which is also referred to as hyperparameter tuning and is commonly employed in ML applications. The most prominent type of hyperparameters (HPs) are those associated with the learning algorithm, which specify its configuration (e.g., the minimum number of observations in each terminal node for tree-based algorithms). If selected by an adequate (and ideally automated) tuning procedure, HPs can substantially enhance the performance of the resulting prediction model. However, HP tuning also complicates model evaluation, as common procedures such as simple $k$-fold cross-validation no longer guarantee an unbiased assessment [78, 104].
@@
An additional challenge comes from the fact that, beyond algorithm HPs, there are also preprocessing HPs, which specify the steps applied to the data before it is fed into the learning algorithm (e.g., selecting the set of features for prediction or determining how missing feature values are handled; [77, 78]). While the tuning of algorithm HPs is rightfully considered important for model performance, the relevance of tuning preprocessing HPs should not be overlooked. Preprocessing steps can make or break a model’s predictive performance, and solely relying on user expertise to specify these steps (which is the alternative to tuning) is often impractical and may result in arbitrary decisions [111]. Despite this, reports of tuning preprocessing HPs aside from feature selection are relatively rare. This could be because integrating preprocessing HPs into automated tuning workflows typically requires advanced programming expertise, which not all applied ML users have, or because this possibility is not widely recognized. Importantly, the limited use of automated tuning procedures for preprocessing HPs does not mean that these HPs are not being tuned at all. In fact, it appears fairly common for applied ML users to experiment informally with different preprocessing options [104, 100, 114], often without realizing that this constitutes a form of (manual) HP tuning. If this type of tuning is indeed conducted subconsciously, it will also remain unaccounted for during model evaluation, thereby increasing the risk of drawing overly optimistic conclusions about the model’s performance.
@@
To avoid such issues, it is essential to educate users in applied settings about the different types of HPs, the different forms of HP tuning, and how tuning can impact both the true and estimated performance of prediction models. Although valuable literature already exists describing the concept of HP tuning and various automated procedures [78, 76, 95, e.g.,], this research primarily adopts the perspective of ML methods researchers who are concerned with evaluating the overall performance of ML algorithms used to generate prediction models. This focus does not align with the perspective of applied ML users, who are more interested in the performance of a specific prediction model. Although this literature is still useful for them—since the general principles described there essentially hold for all types of audiences—applied ML users additionally need specific guidance for developing their “final model” (a notion that does not exist in the methodological context). Moreover, they may find it challenging to extract the relevant insights from literature aimed at a different audience with partly different needs. In contrast, literature explicitly directed toward applied ML users tends to either focus on general guidelines for ML-based predictive modeling, lacking detailed coverage of HP tuning [111, 121, 114, 107, 122, 84, 137, e.g.,], or addresses HP tuning only within specific research areas [104, 91, e.g.,]. Additionally, much of the existing HP tuning literature does not consider preprocessing HPs. Exceptions include the review by [78], which, however, touches on this topic only briefly. This lack of detail is reasonable, given that preprocessing HPs can, in principle, be tuned using the same automated procedures as algorithm HPs. However, this perspective overlooks that preprocessing HPs are often tuned manually in applied settings, which carries implications different from those associated with automated tuning.
@@ 2.1 Terminology and notation
The following terminology and notation is adapted from [78]. Let $\mathcal{D}_{\text{train}}$ be a labeled data set with $n_{\text{train}}$ observations. Accordingly, each observation $i$ ($i=1,\dots,n_{\text{train}}$) consists of an outcome $y^{(i)}$ (i.e. the variable to be predicted, also referred to as label or target) and a $p$-dimensional feature vector $\boldsymbol{x}^{(i)}$ (i.e. the $p$ variables used to predict $y^{(i)}$, also referred to as predictors), where $y^{(i)}$ and $\boldsymbol{x}^{(i)}$ can take any value from the outcome space $\mathcal{Y}$ and feature space $\mathcal{X}$, respectively. Two common types of prediction problems are regression, for which $y^{(i)}$ can be any real number (i.e. $\mathcal{Y}=\mathds{R}$), and classification, for which $y^{(i)}$ can be one of $g$ classes (i.e. $\mathcal{Y}$ is finite and categorical with $|\mathcal{Y}|=g$). We assume that the observations in $\mathcal{D}_{\text{train}}$ are independent and have been sampled from the same (unknown) probability distribution $\mathds{P}_{xy}$.
@@
The general aim of supervised ML is to “learn” a model from the data set $\mathcal{D}_{\text{train}}$ that is able to predict the outcome values of new observations. Essentially, a prediction model is a function $\hat{f}:\mathcal{X}\rightarrow\mathds{R}^{g}$ that maps any observed feature vector $\boldsymbol{x}$ to a prediction vector $\hat{f}(\boldsymbol{x})$ in $\mathds{R}^{g}$. The prediction vector $\hat{f}(\boldsymbol{x})$ either directly corresponds to the predicted outcome value (e.g., for regression, where $g=1$) or can be transformed accordingly (e.g., for classification, where $\hat{f}(\boldsymbol{x})$ corresponds to predicted probabilities for each class and the predicted class could be the class with the highest probability). The prediction model results from a learning pipeline $\mathcal{I}$, which uses the data set $\mathcal{D}_{\text{train}}$ to find the function $\hat{f}$ that yields the best predictions for the true outcome values in $\mathcal{D}_{\text{train}}$. To stress that a prediction model $\hat{f}$ is based on learning pipeline $\mathcal{I}$ and data set $\mathcal{D}_{\text{train}}$, we write $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$. The prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ can usually be parameterized, meaning that it is defined by a set of parameters $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ (simply denoted as $\boldsymbol{\hat{\theta}}$ when data set and learning pipeline are clear from context and $\boldsymbol{\theta}$ when referring to the parameters prior to estimation).
@@
There are two key processes associated with $\mathcal{I}$ and $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$, which we will explore in more detail throughout the paper: (i) the training process, in which the learning pipeline $\mathcal{I}$ is applied to $\mathcal{D}_{\text{train}}$ and estimates the parameters $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ and thus the prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$, and (ii) the prediction process, in which $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ is used to make predictions for an observation (whether from $\mathcal{D}_{\text{train}}$ or from a new data set) with feature vector $\boldsymbol{x}$, resulting in $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}(\boldsymbol{x})$. Note that to make predictions on a new data set, the outcome does not need to be observed (it would only be necessary for evaluating those predictions). The training and prediction processes serve as the foundation for more complex processes related to the development of prediction models, which we will address in Section 2.4.
@@ 2.2 Learning pipeline
Each learning pipeline $\mathcal{I}$ contains a learning algorithm as a central component but can also include several preprocessing steps that are performed before the algorithm is applied to the data. Since preprocessing steps are a particular focus of this paper, we use the term “learning pipeline” instead of the more common term “learner” to emphasize that $\mathcal{I}$ can consist of several components. Note that for now, we consider all components of $\mathcal{I}$ as fixed, but we will discuss the case in which they can be modified in Section 2.3.
@@ 2.2.1 Learning algorithm
The choice of learning algorithm usually depends on the specific prediction problem. For example, if the desired prediction model is a decision tree (which is the case for the real-world prediction problem considered in Section 5), a possible algorithm choice is the well-known Classification and Regression Tree algorithm (CART), which partitions the feature space $\mathcal{X}$ by a sequence of binary splits into terminal nodes and assigns a prediction value to each terminal node [81]. In this case, the parameters of the learning algorithm contained in $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ are the splitting rules that generate the tree structure (i.e. which features are used with which threshold value) and the prediction values at each terminal node. The learning algorithm can also consist of multiple individual algorithms that are combined into one overall algorithm (e.g., random forests). These types of algorithms are referred to as ensemble methods, but will not be discussed further in this paper. In general, the choice of algorithm has a large impact on the hypothesis space of the learning pipeline, i.e. the set of prediction models the learning pipeline can generate. For example, selecting a standard linear regression as algorithm (with $\boldsymbol{\hat{\theta}}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ containing the regression coefficients) would imply that the corresponding learning pipeline would not be able to learn prediction models that do not correspond to linear combinations of the features (e.g., polynomials).
@@ 2.2.2 Preprocessing
While a data set can, in theory, be fed directly into the algorithm (i.e. the algorithm is the only component of the learning pipeline), it typically undergoes some modification first. This process can be referred to as data preprocessing and encompasses all the steps taken to transform the data set from its rawest available form into the final form provided as input to the learning algorithm [107]. Data preprocessing steps are usually performed to improve the performance of the resulting prediction model, to enable the data to be (better) handled by the learning algorithm [135], or to improve the interpretability of the resulting prediction model. To better illustrate the different characteristics of preprocessing steps and their implications on the training and prediction process, we consider a simple learning pipeline as an example, which is also depicted in Figure 1 (middle panel).
@@
Based on this example learning pipeline, we can make a first distinction between preprocessing steps. This distinction concerns whether the steps have parameters estimated from $\mathcal{D}_{\text{train}}$ (with these parameters included in $\boldsymbol{\theta}$) or whether they are parameterless and are carried out independently for each observation [107, 77]. In the example, the replacement of missing feature values is a parameterized preprocessing step, as it involves the parameter $\theta_{impute}$, representing the mean of all non-missing values estimated from $\mathcal{D}_{\text{train}}$. In contrast, the log-transformation of features does not involve any parameters. Other examples of preprocessing steps with parameters include centering or scaling of features, where parameters such as the mean or standard deviation are estimated from $\mathcal{D}_{\text{train}}$. On the other hand, creating a new feature by summing multiple features serves as another example of a parameterless preprocessing step.
@@
We first consider preprocessing steps that affect only the feature distribution. These comprise all preprocessing steps mentioned above, including those in the example learning pipeline. Additional examples are dimensionality reduction techniques (e.g., principal component analysis), feature selection, or data cleaning steps that do not alter the outcome distribution (e.g., correction of errors in features) [111, 135]. Preprocessing steps of this type must be applied not only during training but also during prediction, in the same sequence as in the learning pipeline. This ensures that the model produced by the learning algorithm receives the data in the same format during prediction as it did during training, preserving the validity of the model [77]. This requirement implies that these steps are not only components of the learning pipeline $\mathcal{I}$ but also part of the resulting prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$. Consequently, if a learning pipeline $\mathcal{I}$ includes $h$ preprocessing steps that only affect the feature distribution, the prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}$ is not a single function but a function composition of $h+1$ functions (omitting $\mathcal{D}_{\text{train}}$ and $\mathcal{I}$ for simplicity of notation):
@@
where $\hat{f}_{h+1}$ corresponds to the model resulting from the learning algorithm, and $\hat{f}_{h},\dots,\hat{f}_{1}$ reflect the $h$ preprocessing steps. Accordingly, a more accurate name for a prediction model would be prediction model pipeline, but for brevity, we will continue to use the former. Returning to the example learning pipeline, the resulting prediction model is a composition of three functions, $\hat{f}_{3}(\hat{f}_{2}(\hat{f}_{1}(\boldsymbol{x})))$, where $\hat{f}_{1}$, $\hat{f}_{2}$, and $\hat{f}_{3}$ correspond to the imputation step, the log-transformation step, and the decision tree model, respectively. When making a prediction for one or more observations, all three functions must be applied (see Figure 1, right panel). Importantly, if any functions constituting the prediction model are omitted during the prediction process, or if any preprocessing or algorithm parameters are re-estimated on a new data set for which predictions are to be made, the validity of the prediction model may be compromised. However, in practice, this pitfall is often unavoidable for users who wish to apply a model but were not involved in its development, as studies introducing new prediction models frequently fail to report the preprocessing steps performed prior to applying the learning algorithm [107].
@@
In contrast to preprocessing steps that only affect the feature distribution, preprocessing steps that modify the outcome distribution are not necessarily applied during prediction. Here, we must distinguish between steps aimed at improving compatibility with the learning algorithm and those intended to alter the scope or interpretation of the prediction model. An example of the first type is (invertible) transformations applied to the outcome during training, such as a log-transformation to reduce skewness. To ensure predictions are returned on the correct scale, these transformations must be reversed during prediction [135]. For instance, if the outcome was log-transformed during training, the model will output $\mathrm{log}(\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}}(\boldsymbol{x}))$, which must then be exponentiated to restore the prediction to its original scale. Note that some other compatibility-focused steps are not applied at all during prediction. In the context of classification problems, this includes class-balancing steps such as oversampling, where observations from the least prevalent class are randomly resampled to overcome class imbalance effects during the training process (see, e.g., [111], for more details). In the notation of the prediction model as a function composition introduced above, preprocessing steps that are applied only in their inverted form or not at all during prediction are represented as inversion function or identity function, respectively.
@@
In contrast, preprocessing steps that modify the outcome to alter the scope or interpretation of the prediction model should be consistently applied during prediction. For example, if a continuous outcome is discretized to convert a regression problem into a classification problem [100], this (irreversible) transformation must also be applied to the true outcome during prediction in order to enable a meaningful comparison between the predictions and the actual outcome values. Such transformations of the outcome are not part of the prediction model itself (which maps $\boldsymbol{x}$ to predictions, not $y$), but must be performed alongside the prediction process. Moreover, since the outcome values are generally unknown when making predictions for observations from a new data set that does not correspond to $\mathcal{D}_{\text{train}}$, these transformations are typically not actual steps executed when making predictions but instead determine how the predictions are interpreted.